When working with probabilities why is it sometimes necessary to add and sometimes to multiply?

Whether you need to add or multiply depends on whether it is a case of 'AND' or 'OR'. For 'AND' you multiply, and for 'OR' you add. Below are examples of each case, using dice as an example: To calculate the probability that I get a six on my first roll AND get a five on my second roll, I need to multiply: 1/6 x 1/6 = 1/36To calculate the probability that I get either a six OR a five on my first roll, I need to add: 1/6 + 1/6 = 2/6 = 1/3




MW

Related Maths GCSE answers

All answers ▸

Solve 5x – 2 > 3x + 11


Solve the inequality: 5x-3 < 7x-6


Why do square roots have more than 1 solution?


Factorise and solve x^2 - 8x + 15 = 0